Find two consecutive integers such that the sum of three-fifths of the first and one-fourth of the second is 13.
Answer by mananth(16946)

let the two consecutive numbers be x & x+1
(3/5)x +(1/4)(x+1) =13
LCM of denominator is 20
multiply equation by 20
12x+5(x+1) =20*13
12x+5x+5=260
17x=260-5
17x=255
/17
x=15
the numbers are 15 & 16